OpenStudy (anonymous):

Andy hit 15 home runs this season. If this is 9 fewer than last season how many home runs did he hit last season

OpenStudy (anonymous):

6

OpenStudy (anonymous):

nope...24...srry

OpenStudy (anonymous):

What is the formula I write

OpenStudy (lifeisadangerousgame):

Well, if he hit nine less than last season, then we need to add 9 to 15. The formula would be...x - 9 = 15 I think

OpenStudy (charlotte123):

15 + 9 = x?

OpenStudy (anonymous):

15=x-9

OpenStudy (charlotte123):

Or x - 9 = 15 ^_^

OpenStudy (charlotte123):

either way :3

OpenStudy (anonymous):

THANKS!!

OpenStudy (charlotte123):

np :D GREAT JOB EVERYONE ^_^